OpenStudy (anonymous):

solve using matrices 3x-y+2z=6 -x+y=2 x-2z=-5

OpenStudy (anonymous):

Here's that problem expressed in matrices. If you can solve it, go for it! Otherwise, drop me a line and I'll explain further. \[\left[\begin{matrix}3 & -1 & 2 \\ -1 & 1 & 0 \\ 1 & 0 & -2\end{matrix}\right] \left(\begin{matrix}x \\ y \\ z\end{matrix}\right)=\left(\begin{matrix}6 \\ 2 \\ -5\end{matrix}\right)\] Here's a link to the MIT Linear Algebra group if you're going to have more questions on this subject: http://openstudy.com/study?version=feed:join-study-group&referrer=mit%2018.06%20linear%20algebra,%20spring%202010&domain=ocw.mit.edu#/groups/MIT%2018.06%20Linear%20Algebra%2C%20Spring%202010